### v3

# A Panchromatic View of Late-time Shock Power in the Type II Supernova 2023ixf

Abstract
We present multi-wavelength observations of the type II supernova (SN II) 2023ixf during its first two years of evolution. We combine ground-based optical/NIR spectroscopy with Hubble Space Telescope (HST) far- and near-ultraviolet spectroscopy and James Webb Space Telescope (JWST) near- and mid-infrared photometry and spectroscopy to create spectral energy distributions of SN 2023ixf at +374 and +620 days post-explosion, covering a wavelength range of m. The multi-band light curve of SN 2023ixf follows a standard radioactive decay decline rate after the plateau until days, at which point shock powered emission from ongoing interaction between the SN ejecta and circumstellar material (CSM) begins to dominate. This evolution is temporally consistent with 0.3-10 keV X-ray detections of SN 2023ixf and broad “boxy” spectral line emission, which we interpret to signal reprocessing of shock luminosity in a cold dense shell located between forward and reverse shocks. Using the expected absorbed radioactive decay power and the detected X-ray luminosity, we quantify the total shock powered emission at the +374 and +620 day epochs and find that it can be explained by nearly complete thermalization of the reverse shock luminosity as SN 2023ixf interacts with a continuous, “wind-like” CSM with a progenitor mass-loss rate of M⊙ yr-1 (). Additionally, we construct multi-epoch spectral models from the non-LTE radiative transfer code CMFGEN, which contain radioactive decay and shock powers, as well as dust absorption, scattering, and emission. We find that models with shock powers of erg s-1 and of silicate dust in the cold dense shell and/or inner SN ejecta can effectively reproduce the global properties of the late-time ( days) UV-to-IR spectra of SN 2023ixf.
1 Introduction
Supernova (SN) 2023ixf is a type II supernova (SN II) that occurred on 2023-05-18 in Messier 101 at a distance of Mpc (Riess et al., 2022). SN 2023ixf is one of the closest SN in the last decade and, consequently, was reported to the Transient Name Server (TNS) within 1 day of first light by Koichi Itagaki (Itagaki, 2023) on 2023 May 19 17:27:15 (MJD 60083.73). The earliest reported detection of SN 2023ixf was on MJD 60082.85 at mag (-12.1 mag) in -band following a deep upper limit of mag ( mag) on MJD 60082.66 (Mao et al., 2023; Li et al., 2024). Consequently, an estimate of the time of first light is MJD , which is adopted throughout this work.
The early-time spectra of SN2023ixf exhibit emission lines from H i, He i/ii, C iv, and N iii/iv/v, superimposed on a hot, blue spectral continuum (Jacobson-Galán et al., 2023; Bostroem et al., 2023; Teja et al., 2023; Zhang et al., 2023). These lines, commonly referred to as “flash” or “IIn-like” features, likely originate from the sustained photoionization of dense circumstellar material (CSM) located ahead of the forward shock (Gal-Yam et al., 2014; Khazov et al., 2016; Yaron et al., 2017; Dessart et al., 2017; Terreran et al., 2022; Bruch et al., 2021, 2023; Dessart & Jacobson-Galán, 2023; Jacobson-Galán et al., 2024a). Much like Type IIn supernovae, the emission line profiles consist of a narrow core accompanied by broad Lorentzian “wings” that extend up to 1000 , a signature of electron scattering within the ionized, optically thick CSM (Chugai, 2001; Dessart et al., 2009, 2017; Huang & Chevalier, 2018). Additionally, SN 2023ixf has the earliest and highest cadence UV spectroscopy (Teja et al., 2023; Zimmerman et al., 2024), spectropolarimetry (Vasylyev et al., 2023; Singh et al., 2024; Shrestha et al., 2025; Vasylyev et al., 2025) and high resolution spectroscopy (Smith et al., 2023; Zimmerman et al., 2024; Dickinson et al., 2025) of any CSM-interacting SN II to date.
Multi-band photometry of SN 2023ixf began within hours of first light, capturing shock emergence from a dense, compact CSM at cm. The early light curve shows a two-component rise: the first 12 hours require a distinct power-law fit, while the later evolution (h) follows a typical trend (Hosseinzadeh et al., 2023). During its first day, SN 2023ixf exhibited a rapid red-to-blue color evolution, common among CSM-interacting SNe II with IIn-like features (e.g., SN 2024ggi; Jacobson-Galán et al. 2024b; Shrestha et al. 2024), likely marking shock breakout from the innermost CSM. This coincided with a nearly constant blackbody radius and rising temperature (Zheng et al., 2025; Zimmerman et al., 2024). Li et al. (2024) attributes the early red colors to dust destruction during breakout, consistent with the IR excess detected by NEOWISE-R at days (Van Dyk et al., 2024a).
The dust-enshrouded red supergiant (RSG) progenitor star of SN 2023ixf was clearly detected in multiple HST optical bands, in Channels 1 and 2 of Spitzer, and in and bands from Gemini-North and MMT. Estimates of the progenitor’s Zero Age Main Sequence (ZAMS) mass vary widely, ranging from 8–20 (Kilpatrick et al., 2023; Jencson et al., 2023; Niu et al., 2023; Soraisam et al., 2023; Van Dyk et al., 2024b; Qin et al., 2023; Pledger & Shara, 2023; Xiang et al., 2024; Ransome et al., 2024) and no evidence of a binary companion was found, though only secondary stars with masses above could be excluded. Additional progenitor mass constraints were made from model matching to the SN 2023ixf light curve (Moriya & Singh, 2024; Singh et al., 2024; Bersten et al., 2024; Hsu et al., 2024; Forde & Goldberg, 2025; Cosentino et al., 2025) and nebular spectroscopy (Ferrari et al., 2024; Fang et al., 2025; Kumar et al., 2025; Michel et al., 2025; Li et al., 2025; Folatelli et al., 2025; Jacobson-Galán et al., 2025). Notably, Spitzer data showed long-term variability with a pulsation period of 1000 days (Kilpatrick et al., 2023; Jencson et al., 2023; Soraisam et al., 2023). Given the early-time CSM interaction observed in SN 2023ixf, several studies searched for pre-explosion outbursts, common in some CSM-interacting SNe II (e.g., Strotjohann et al. 2021; Jacobson-Galán et al. 2022), but found no precursor activity in archival X-ray/UV/optical/NIR data (Panjkov et al., 2024; Neustadt et al., 2024; Ransome et al., 2024; Dong et al., 2023; Rest et al., 2025; Flinner et al., 2023).
X-ray observations of SN2023ixf began with Swift-XRT ( day), followed by NuSTAR (3–79 keV; d), Chandra (0.5–8 keV; days), and XMM-Newton (0.3–10 keV; days). This unprecedented soft and hard X-ray coverage enabled the first precise temperature constraints for a SN II with IIn-like features (Grefenstette et al., 2023; Chandra et al., 2024; Panjkov et al., 2024; Nayana et al., 2025). The X-ray emission, dominated by the forward shock, is best fit by an absorbed bremsstrahlung model (Nayana et al., 2025; Chandra et al., 2024; Panjkov et al., 2024), peaking at erg s-1 within a week. Detection of Fe K and high intrinsic confirms early X-ray absorption by dense, confined CSM (Grefenstette et al., 2023; Nayana et al., 2025) and the observed rising in X-ray flux reflects decreasing photoelectric absorption as CSM density also decreases and the ionization fraction potentially rises. A high-cadence radio campaign began at d with SMA at mm wavelengths, though no emission was detected until d with the VLA at 10 GHz (Berger et al., 2023; Matthews et al., 2023). Additional broad-band radio spectral coverage was provided by GMRT, LOFAR, NOEMA, the Japanese and Korean VLBI Networks, and EVN (Nayana et al., 2025; Timmerman et al., 2024; Iwata et al., 2025; Lee et al., 2024). The radio emission, produced by synchrotron radiation from the forward shock, is initially mostly suppressed by free–free absorption, which weakens as the CSM density decreases at larger shock radii (Nayana et al., 2025). Despite potentially contradictory mass-loss rate estimates between optical spectroscopy and X-ray observations within the first week post-explosion, long-term X-ray and radio monitoring provides a consistent mass-loss rate of M⊙ yr-1 at shock radii cm (Nayana et al., 2025; Panjkov et al., 2024). A complete review of all observations and analysis of SN 2023ixf to date is presented in Jacobson-Galán (2025).
In this paper, we present multi-wavelength observations and analysis of the continuous 0.1-30 m SN 2023ixf SED at late-time phases (Fig. 1). In §2, we present observations and reduction techniques. In §3, we present an analysis of the radioactive decay and shock powered emission in the SN 2023ixf SED. Conclusions are then presented in §4. All phases reported in this paper are with respect to the adopted time of first light () and are in rest-frame days, as denoted by . We adopt a redshift of (Perley et al., 2023), and Milky Way and host galaxy reddening values of mag and mag (Jacobson-Galán et al., 2023).
2 Observations
2.1 Photometric Observations
SN 2023ixf was observed with the Pan-STARRS telescope (PS1/2; Kaiser et al., 2002; Chambers et al., 2017) between 2023-05-21 and 2025-06-02 in -bands through the Young Supernova Experiment (YSE) (Jones et al., 2021). Data storage/visualization and follow-up coordination was done through the YSE-PZ web broker (Coulter et al., 2022, 2023). The YSE photometric pipeline is based on photpipe (Rest et al., 2005), which relies on calibrations from (Magnier et al., 2020; Waters et al., 2020). Each image template was taken from stacked PS1 exposures, with most of the input data from the PS1 3 survey. All images and templates were resampled and astrometrically aligned to match a skycell in the PS1 sky tessellation. An image zero-point is determined by comparing PSF photometry of the stars to updated stellar catalogs of PS1 observations (Flewelling et al., 2016). The PS1 templates are convolved with a three-Gaussian kernel to match the PSF of the nightly images, and the convolved templates are subtracted from the nightly images with HOTPANTS (Becker, 2015). Finally, a flux-weighted centroid is found for the position of the SN in each image and PSF photometry is performed using “forced photometry”: the centroid of the PSF is forced to be at the SN position. The nightly zero-point is applied to the photometry to determine the brightness of the SN for that epoch.
We obtained imaging of SN 2023ixf with the Las Cumbres Observatory (LCO) 1 m telescopes from 2023-05-20 to 2024-09-24 (Programs NSF2023A-011, NSF2023A-015, NSF2023B-004, ANU2024A-004, ANU2024B-003, and ANU2025A-003; PIs Foley, Kilpatrick, Auchettl). After downloading the BANZAI-reduced images from the LCO data archive (McCully et al., 2018), we used photpipe (Rest et al., 2005) to perform DoPhot PSF photometry (Schechter et al., 1993). All photometry was calibrated using PS1 stellar catalogs described above with additional transformations to SDSS -band derived from Finkbeiner et al. (2016). For additional details on our reductions, see Kilpatrick et al. (2018). We also observed SN 2023ixf with the Lulin 1 m and Thacher (Swift et al., 2022) telescopes in bands from 2023-05-21 to 2025-04-20. Standard calibrations for bias and flat-fielding were performed on the images using IRAF, and we reduced the calibrated frames in photpipe using the same methods described above for the LCO images. We obtained NIR photometry using the Wide-field InfraRed Camera (WIRC; Wilson et al. 2003) on the 200-inch telescope at Palomar Observatory. The WIRC data were reduced using standard methods for dark subtraction, flat-fielding, sky subtraction, astrometric and photometric calibration (De et al., 2021).
The Ultraviolet Optical Telescope (UVOT; Roming et al. 2005) onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004) observed SN 2023ixf from 2023-05-19 to 2025-02-23 ( days). We supplement the published early-time UVOT photometry from Zimmerman et al. (2024) with reductions at days. We performed aperture photometry with a 5′′ region radius with uvotsource within HEAsoft v6.33 (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc)111We used the most recent calibration database (CALDB) version., following the standard guidelines from Brown et al. (2014)222https://github.com/gterreran/Swift_host_subtraction. In order to remove contamination from the host galaxy, we employed pre-explosion images to subtract the measured count rate at the location of the SN from the count rates in the SN images and corrected for point-spread-function (PSF) losses following the prescriptions of Brown et al. (2014).
In addition to our observations, we include -band photometry from the Zwicky Transient Facility (ZTF; Bellm et al. 2019; Graham et al. 2019) forced-photometry service (Masci et al., 2019). We also include multi-band photometry presented in Jacobson-Galán et al. (2023); Van Dyk et al. (2024a); Hsu et al. (2024); Zimmerman et al. (2024); Singh et al. (2024) in our analysis. Furthermore, we include JWST MIRI imaging of SN 2023ixf from 2024-03-15 (JWST-GO-3921, PI: Fox) and 2025-01-27 (JWST-GO-5290, PI: Ashall), which includes observations in F770W, F1000W, F1500W, F1800W, F2100W, and F2500W filters. We performed PSF photometry on all filters and epochs using the spacephot photometry package and the PSF models from WebbPSF.333https://space-phot.readthedocs.io/en/latest/ All photometric measurements are consistent with those derived in Medler et al. (2025). The complete UVOIR light curve of SN 2023ixf is shown in Figure 2.
2.2 Spectroscopic Observations
We obtained late-time spectra of SN 2023ixf with the Kast spectrograph on the 3-m Shane telescope at Lick Observatory (Miller & Stone, 1993). For all of these spectroscopic observations, standard CCD processing and spectrum extraction were accomplished with IRAF444https://github.com/msiebert1/UCSC_spectral_pipeline. The data were extracted using the optimal algorithm of Horne (1986). Low-order polynomial fits to calibration-lamp spectra were used to establish the wavelength scale and small adjustments derived from night-sky lines in the object frames were applied. Additional optical spectra were obtained with Keck/LRIS (Oke et al., 1995) and reduced with Lpipe (Perley, 2019). Near-infrared (NIR) spectra were obtained with the Near-Infrared Echelle Spectrograph (NIRES; Wilson et al. 2004) located on Keck-II. The data were reduced using a custom version of the IDL based reduction package Spextool (Cushing et al., 2004) modified for use with NIRES as well as the Pypeit spectral reduction pipeline (Prochaska et al., 2020). For the Spextool reduction, we used xtellcor (Vacca et al., 2003) to correct for telluric features in our spectrum using an A0 standard star observed close in airmass and time to our target. NIRES data from days were obtained through the Keck Infrared Transient Survey (KITS; Tinyanont et al. 2024).
We observed SN 2023ixf on 2024-12-31 starting at 15:44 UT with both the Medium and Small slicers of the Keck Cosmic Web Imager (KCWI; blue channel, Morrissey et al. 2018) and the Keck Cosmic Reionization Mapper (KCRM; red channel) mounted on the Keck II telescope. For the Medium slicer observations, the blue channel was configured with the BL grating centered at 4500 Å and the red channel used the RL centered at 7150 Å. We acquired one 850s image in the blue channel and 2x300s in the red channel, both channels configured with the 2x2 binning. For the Small slicer, we used the BL grating centered at 4500 Å and a 2x2 binning to acquire 2x650s images, while for the red channel we used the RH1 grating () centered at 6520 Å and a 1x1 binning to target the H line of SN 2023ixf.
Far- and near-UV spectra of SN 2023ixf were obtained on 2024-03-24 ( days) and 2025-01-26 ( days) with the Hubble Space Telescope (HST) using the Multi-Anode MicroChannel Array (MAMA) detectors on the Space Telescope Imaging Spectrograph (STIS) through HST programs 17497 (PI Valenti) and 17772 (PI Bostroem) Bostroem et al. (2024a, 2025). Standard bias subtraction, flat-fielding and cosmic ray rejection were performed automatically before downloading from the Mikulski Archive for Space Telescopes (MAST). Wavelength calibration, flux calibration and 1-D extraction was performed using the calstis pipeline routine. Near- and mid-infrared spectroscopy of SN 2023ixf were obtained on 2024-01-26, 2024-05-26, and 2025-01-08 ( days) (JWST-DD-4575, JWST-GO-5290; PI: Ashall) with the James Webb Space Telescope (JWST) using NIRSpec with G235M-F170LP and G395M-F290LP grating/filter combinations (Jakobsen et al., 2022) and MIRI/LRS (Kendrew et al., 2015). These data were reduced using the jwst555https://github.com/spacetelescope/jwst pipeline, which performs standard bias subtraction, flat-fielding, wavelength and flux calibrations, and spectral extraction. The complete optical/NIR spectral sequence is shown in Figure 3 and a log of all spectral observations utilized in this paper is provided in Table A1. All photometric and spectroscopic data will be made public online.666https://github.com/wynnjacobson-galan/SN2023ixf
2.3 X-ray Observations with Swift-XRT
The X-Ray Telescope (XRT, Burrows et al. 2005) on board the Swift spacecraft (Gehrels et al., 2004) observed the field of SN 2023ixf from 2023-05-19 to 2025-02-23 ( days). Observations across multiple X-ray telescopes, including XRT during the first 200 days of SN 2023ixf have been analyzed and published in recent studies (e.g., Grefenstette et al. 2023; Panjkov et al. 2024; Chandra et al. 2024; Zimmerman et al. 2024; Nayana et al. 2025). We focus on the XRT epochs from days, which cover the range of late-time UVOIR observations discussed in §2.1 & §2.2. We analyzed the data using HEAsoft v6.33 and followed the prescriptions detailed in Margutti et al. (2013), applying standard filtering and screening. Below we outline our first method for modeling and analyzing the XRT data with Xspec, but for completeness we also outline a secondary modeling method in Appendix §A. For this first method, we chose to merge the event files and use the combined epoch for analysis of the X-ray spectrum in the following date ranges: 2023-11-21 to 2023-12-05 ( ks), 2024-09-29 to 2024-10–24 ( ks), and 2024-12-23 to 2024-12-24 ( ks). In all merged images, a bright source of X-ray emission is clearly detected at the SN location with significance of against the background.
From each merged event file, we extracted a spectrum using a 25′′ region centered at the location of SN 2023ixf and correct for background emission with a 100′′ source-free region. Similar to Grefenstette et al. (2023), Chandra et al. (2024), Panjkov et al. (2024), and Nayana et al. (2025), we use Xspec to model each 0.3-10 keV spectra with an absorbed thermal bremsstrahlung model (tbabs*ztbabs*bremss), which includes solar abundances (Asplund et al., 2009), and a line-of-sight hydrogen column density of cm-2. Furthermore, each model spectrum includes an additional power law component in order to account for the contaminating source presented in Nayana et al. (2025). We fit each X-ray spectrum with two separate models: (1) two temperature components for the forward shock (FS) and reverse shock (RS) and (2) single temperature model for only the FS. We fix the temperatures in each spectral fit so that the FS and RS temperature evolves as (Nayana et al., 2025) and , assuming an ejecta density profile power-law index of (Chevalier & Fransson, 2017; Chandra et al., 2024). As shown in Appendix Figure A.1, the relatively small number of counts and the lower Swift-XRT sensitivity at 1 keV inhibits a meaningful constraint on intrinsic ahead of the FS, which is likely cm-2 based on the measured decline rate in Nayana et al. (2025). When fitting with a RS model component, we invoke a time-dependent CDS column density of cm-2 assuming a “wind-like” CSM density profile (e.g., see Chevalier & Fransson 2017; Nayana et al. 2025), which is fixed in each model fit. To determine whether an additional temperature component for the RS is necessary, we calculate the Bayesian Information Criterion (BIC) for each model fit to the SN 2023ixf X-ray spectrum. As shown in Table A2, there is no statistical evidence that a RS emission component is needed to best fit the X-ray data. Furthermore, we report the 3 uncertainties on the normalization parameter for the RS component and find that the best-fit value is consistent with zero for the first two epochs and unphysical for the third epoch. All X-ray modeling parameters from this method and that described in §A are reported in Table A2.
In Figure 4 we present the unabsorbed 0.3 - 10 keV X-ray light curve from Nayana et al. (2025) and the additional luminosities derived from our analysis using a single temperature model fit. While the day data are consistent with the decline rate expected from pure FS emission, the flattening of the keV X-ray light curve at days may be indicative of emerging RS emission and/or increasing CSM density. To quantify the CSM density at shock radii cm, we use the normalization of the X-ray spectrum to calculate the total emission measure (EM) at each epoch, which can then be converted to unshocked CSM density (e.g., see Eqn. 2 in Brethauer et al. 2022). We adopt all the same parameters as Nayana et al. (2025) when applying this EM formalism. As shown in Figure 4, the CSM density continues to decline as , which can be matched by a steady-state mass loss rate of M⊙ yr-1 (). Overall, we find no evidence for a RS contribution to the observed keV X-ray luminosity shown in Figure 4.
3 Analysis
3.1 Radioactive Decay Power
As shown in Figure 2, the photospheric phase persists as a linearly-declining plateau in the SN 2023ixf light curve for 83 days. Following the fall from the plateau, the photometric evolution of SN 2023ixf becomes consistent with radioactive decay from the standard Fe chain (Arnett, 1982). Modeling of the nebular decline rate has suggested and incomplete -ray trapping with a characteristic timescale of days (Zimmerman et al., 2024; Singh et al., 2024; Li et al., 2025). To confirm these parameters, we construct two variations of the post-plateau bolometric light curve, which are then fit with analytic formalisms for radioactive decay power. First, we construct a pseudobolometric light curve using filters and fit this light curve with the pseudobolometric light curve of SN 1987A derived with the same filter combination (e.g., see Eqn. 1 of Jacobson-Galán et al. 2025), which yields and days. We note that consistent values are found when this method is applied to the more confined filter combination. Second, we construct a pseudobolometric light curve from the combination of filters and which covers m. Given the emergence of shock powered emission at late-times (§3.2), we model this pseudobolometric light curve at days and find and days. We chose to exclude UV and mid-IR photometry in this light curve given that (1) of SN II luminosity is expected to be emitted between m at days (Dessart, 2025a) if it arises from decay power only and (2) the UV excess from persistent shock powered emission shown in Figure 5 would lead to an overestimate of . Based on the average of both modeling attempts, we choose to adopt and days when calculating the contribution of absorbed radioactive decay power to the observed luminosity of SN 2023ixf.
3.2 Shock Powered Emission
As shown in Figures 2 & 5, an excess emission beyond what is expected from pure radioactive decay first becomes apparent at UV wavelengths as the Swift-UVOT light curve becomes flat at days and remains bright at mag above the host galaxy background – these observations are also consistent with observed emission in the near- and far-UV at days (Bostroem et al., 2025). Host galaxy subtracted optical and NIR photometry shows a similar flattening at days and the pseudobolometric light curve (e.g., Fig. 5) deviates from a radioactive decay decline rate as shock powered emission from persistent SN ejecta-CSM interaction becomes the dominant power source in SN 2023ixf. Here we define “shock powered emission” as all additional UVOIR luminosity that does not come from absorbed radioactive decay power e.g., the reprocessed X-ray luminosity from the radiative RS. The presence of shock powered emission in SN 2023ixf is also confirmed by the spectral evolution shown in Figure 3 as the forbidden emission lines observed at year fade and prominent boxy H, H, P, He i and [Fe ii] emission become dominant at days – this transition in H is shown in Figure 6. The presence of these emission line profiles is direct evidence for a radiative reverse shock that injects power into both the CDS and the inner ejecta (Nymark et al., 2006; Chevalier & Fransson, 2017; Dessart & Hillier, 2022). Similarly, prominent Ly and Mg ii emission is observed at UV wavelengths, likely from the CDS (e.g., Fransson et al. 2005; Dessart et al. 2023; Bostroem et al. 2024a, 2025. Furthermore, evidence for ongoing CSM interaction is given by persistent X-ray and radio emission, both of which imply a mass loss rate of M⊙ yr-1 at shock radii cm (Panjkov et al., 2024; Chandra et al., 2024; Nayana et al., 2025). However, this shock powered emission source from ejecta interaction with a likely steady-state wind has remained subdominant compared to radioactive decay power until days in optical/IR bands and its emergence at these late times is not the result of SN ejecta interaction with some detached shell of CSM.
In order to quantify the relative contributions from shock and radioactive decay power, we construct SEDs of SN 2023ixf at days, which cover a wavelength range of m. Because the UV spectrum was obtained days, we scale the spectrum based on a linear interpolation of the integrated UV flux at the day and day HST observation epochs. At days, the bolometric luminosity in this wavelength range is erg s-1 and the expected luminosity from radioactive decay power absorbed by the ejecta at this phase is erg s-1 using the parameters derived in §3.1. Consequently, the emergent shock powered emission across this wavelength range is erg s-1 at days. Then, at days, we find erg s-1, erg s-1, and erg s-1. Additionally, within this phase range, the observed X-ray emission (e.g., Fig. 4) remains relatively constant with an unabsorbed 0.3-10 keV luminosity of erg s-1.
As discussed in §2.3, the observed late-time X-ray emission is arising primarily from the FS, which is predicted to be almost completely adiabatic at this phase (Fransson et al., 1996; Chevalier & Fransson, 2017). The RS X-ray emission is completely thermalized in the CDS and re-emitted as the observed shock powered emission in SN 2023ixf. For ejecta-CSM interaction with a constant mass-loss rate, the thermal bremsstrahlung (“free-free”) bolometric emission from an adiabatic shock can be described as:
where is mass-loss rate in units of M⊙ yr-1, is the wind velocity in units of , is the Gaunt factor (assumed to be of order unity), is electron temperature, is ion temperature, for the FS and for the RS (Chevalier & Fransson, 2003, 2017). We choose a wind velocity of , which has uncertainties that account for the first narrow line velocity measurement of at 1.5 days (Dickinson et al., 2025) and any prior radiative acceleration during shock breakout (Dessart, 2025b). We adopt based on the observed X-ray spectrum temperature and the expected ion temperature given a shock velocity of as well as , and M⊙ yr-1. Using these parameters, we find expected forward shock luminosities of erg s-1 at days and erg s-1 at days.
For a radiative RS, we use Eqn 3.17 from Fransson et al. (1996) where the bolometric luminosity goes as:
where is the ejecta velocity at the RS in units of and is the CSM density power-law index. For a radiative cooling RS and a density profile, the time variable above disappears and the luminosity decreases as , which then goes as for and given the time dependence of the ejecta velocity (e.g., see Eqn. 2.2 in Fransson et al. 1996). We adopt M⊙ yr-1() and choose an ejecta velocity of , which covers the maximum velocity observed in the “boxy” H profile (e.g., Fig. 6). Applying Eqn. 3.14 in Fransson et al. (1996), we confirm that the RS is indeed radiative because the cooling time is much less than the age of the SN at both epochs e.g., days at days and days at days. Using these parameters and the equation above, we find erg s-1 at days and erg s-1 at days. At these phases, the shock powered emission () observed in SN 2023ixf is erg s-1 and erg s-1, both of which are consistent with the predicted RS emission. Overall, the emergent shock powered emission observed in SN 2023ixf indicates nearly complete thermalization of RS X-rays converted to UVOIR emission by the CDS and inner ejecta. Intriguingly, the power budget in SN 2023ixf at days implies that the luminous near- and mid-IR emission from silicate and carbonaceous dust grains (e.g., Figures 1 & 8) is a product of shock powered emission since radioactive decay power is insufficient to heat the dust formed in the CDS and inner ejecta.
3.3 Radiative Transfer Model of the Low-Energy Radiation
The multi-epoch, late-time SED of SN 2023ixf enables an unprecedented opportunity to model the UV-to-IR emission of a CSM-interacting SN II. To do this, we compute late-time synthetic spectra with the non-LTE radiative transfer code CMFGEN (Hillier & Dessart, 2012) and employ a combination of SN II nebular models from Dessart et al. (2021), SN II late-time shock power models from Dessart & Hillier (2022); Dessart et al. (2023), and SN II models in which the radiative transfer treats dust absorption, scattering and emission (Dessart et al., 2025). Unlike models that focus on blue/red asymmetry within individual emission lines (e.g., H), these models account for dust absorption and thermal emission across the complete SED, but require dust location/temperature as input parameters in the CMFGEN calculation (i.e., not computed ab initio; see Dessart et al. 2025 for details). We adopt a progenitor ZAMS mass of , which contains a similar 56Ni mass to SN 2023ixf. However, in order to account for the greater -ray escape observed in SN 2023ixf and to match the overall continuum level at late-times, we scale the absorbed radioactive decay power to 40% of its original model value. We explore a range of shock powers ( erg s-1), dust masses (), dust temperatures ( K), dust compositions (silicate vs carbonaceous), dust grain size (m), and dust location (inner ejecta at or within the narrow CDS at ) – a subset of this model exploration is presented in Appendix Figure A.3. We find that of 1m silicate dust with K yields a satisfactory match to the IR spectrum of SN 2023ixf, which shows clear emission from silicate dust at m. We note that the CMFGEN model includes gas (atom and ions) and dust, but ignores molecules.
| Phase | ||||||
| (days) | (erg s-1) | (erg s-1) | (erg s-1) | (erg s-1) | (erg s-1) | (erg s-1) |
| 374 | ||||||
| 620 |
As shown in Figures 7 and A.3, the inclusion of dust can also produce similar blue/red asymmetries that are observed in both the boxy and narrow components of H, indicative of dust formation in the inner ejecta during the first complete spectral epoch ( day) as well as in the CDS by the next epoch ( days). Furthermore, the asymmetries in the line profiles may also be wavelength dependent (e.g., see Fig. 6), which can be used to trace the dust opacity as has been done for type IIn SNe (e.g., SN 2010jl; Gall et al. 2014) and could also suggest larger grain sizes of m. While some pre-existing dust external to the FS may exist, the overall CSM dust mass must be quite small because it would cause significant attenuation at UV wavelengths, which is not observed in the HST spectra. This overall evolution of dust formation is supported by the detection of molecular CO emission beginning at days (Park et al., 2025; DerKacy et al., 2025) and mid-infrared silicate emission (Medler et al., 2025). As shown in Appendix Figure A.2, we fit both epochs of IR spectra with analytic formalisms for dust emission (e.g., see Fox et al. 2010, 2011; Tinyanont et al. 2019; Shahbandeh et al. 2023; Pearson et al. 2025; Tinyanont et al. 2025)777https://github.com/stinyanont/sed_et_al after masking out all emission lines and find consistent silicate dust masses to what is adopted in the best-matched CMFGEN spectra. However, we find that the IR spectrum of SN 2023ixf requires an additional C-rich dust components to match the NIR emission, which is not included in the best-matched CMFGEN model spectra with Si-rich dust. Furthermore, it is clear from the analytic model fits that the emission at m is not well-matched – this could be due to a number of effects e.g., dust opacities, clumping, etc. (Sarangi, 2022; Dessart et al., 2025). Future simulations will explore combinations of various dust compositions, temperatures and structures.
In Figure 8, we present the best-matched CMFGEN models at days compared to the UVOIR SED of SN 2023ixf at days. We find that a shock power of erg s-1 is sufficient to match the total UV emission and the most prominent line profiles such as Ly and Mg ii as well as the boxy emission observed in H at days. At days, we find that the model shock power needs to be reduced to erg s-1 so as to not over-predict the UV emission lines, but also observe that the erg s-1 shock power model is a better match to the optical and NIR spectra at this phase. While the reduction in total UV flux between epochs is expected from a declining RS X-ray luminosity (e.g., Eqn. 1), we note that the observed shock powered emission is larger than the model prediction at days. Furthermore, at days, one CMFGEN model includes dust in the CDS and another with dust in the inner ejecta (, both of which have the most dramatic effect on the H and He emission line profiles. However, these current CMFGEN models show inconsistency with the spectral shape and overall flux in the IR, which is likely the product of a fixed dust temperature and single dust emission location within the SN. Further analysis including hybrid dust locations and dust temperature profiles should be explored but is beyond the scope of this paper.
4 Conclusions
In this paper, we present late-time, multi-wavelength observations of the CSM-interacting SN II 2023ixf spanning X-ray, UV, optical, IR wavelengths. Below we summarize the primary observational findings from this work.
Following the end of the light curve plateau, the multi-band light curve of SN 2023ixf follows a decline rate dictated by absorbed radioactive decay power. Then, at days for UV bands and days for optical/IR bands, the SN 2023ixf light curve flattens as shock powered emission () from enduring CSM-interaction dominates over absorbed radioactive decay power (). We model the band pseudobolometric light curve at days and find , days.
At late-time phases, H shows broad, “boxy” emission from a CDS with velocity of – this structure also being seen in UV (e.g., Mg ii , Ly-) and NIR (e.g., He i, Pa-) emission lines. We interpret these observations as a transition to a regime dominated by shock powered emission where the UV/optical/NIR spectra are dominated by emission from the CDS and inner ejecta. Significant blue-red asymmetry observed in H, combined with luminous near- and mid-IR emission, indicates dust formation in the inner ejecta beginning at days and then in the CDS by at least days.
We quantify the relative contributions from absorbed radioactive decay power and shock powered emission in SN 2023ixf by constructing m SEDs at days. In the first epoch, we find erg s-1 and erg s-1, which results in an emergent shock powered emission of erg s-1. Then, at days, we find erg s-1, erg s-1, and erg s-1.
We estimate unabsorbed 0.3-10 keV luminosities at days by modeling the Swift-XRT spectra with an absorbed thermal bremsstrahlung model. Despite indirect evidence for the presence of a radiative RS from UV-to-IR spectra and the late-time flattening of the X-ray light curve, we find no statistical evidence for a secondary, lower temperature component from this shock in the X-ray spectra. We use the emission measure from the adiabatic FS component to calculate unshocked CSM densities at shock radii of cm and confirm that the CSM density profile continues to be well-described by a constant mass loss rate of M⊙ yr-1 ().
We find that the measured shock powered emission contribution to the m SED at days can be explained by nearly complete thermalization of power from a radiative RS into both the CDS and inner SN ejecta. Using the analytic formalism for thermal bremsstrahlung emission from ejecta-CSM interaction, the observed shock powered emission in SN 2023ixf can be reproduced by RS emission from a constant mass-loss rate of M⊙ yr-1, assuming , and (Eqn. 2). Furthermore, these same parameters predict a FS luminosity that is consistent with the observed X-ray light curve.
We construct CMFGEN model spectra at days, which include absorbed radioactive decay power, shock power and dust emission. As shown in Figure 1, model spectra with erg s-1 and of silicate dust can reproduce the global properties of the late-time SN 2023ixf m SEDs.
Given its present evolution, SN 2023ixf represents an ideal laboratory to study shock powered emission through continued, multi-wavelength observations. Future X-ray spectra should be able to confirm multiple temperature components from the emerging RS and also constrain the CSM densities at shock radii cm. These observations, coupled with complete UVOIR SED coverage will enable more robust constraints on the thermalization efficiency of the RS luminosity and the mechanisms for heating newly formed dust.
5 Acknowledgments
W.J.-G. is supported by NASA through Hubble Fellowship grant HSTHF2-51558.001-A awarded by the Space Telescope Science Institute, which is operated for NASA by the Association of Universities for Research in Astronomy, Inc., under contract NAS5-26555. This work was granted access to the HPC resources of TGCC under the allocation 2024 – A0170410554 made by GENCI, France. C.D.K. gratefully acknowledges support from the NSF through AST-2432037, the HST Guest Observer Program through HST-SNAP-17070 and HST-GO-17706, and from JWST Archival Research through JWST-AR-6241 and JWST-AR-5441. Parts of this research were supported by the Australian Research Council Centre of Excellence for Gravitational Wave Discovery (OzGrav), through project number CE230100016. The Margutti team at UC Berkeley is partially funded by the Heising-Simons Foundation under grants #2018-0911 and #2021-3248 (PI R. Margutti). The TReX team at UC Berkeley is partially supported by the NSF under grant AST-2224255, and by the Heising-Simons Foundation under grant #2021-3248 (PI R. Margutti). V.V.D.’s X-ray astronomy research is currently supported by Chandra Grants G03-24041X and TM4-25004X. K.A.B. is supported by an LSSTC Catalyst Fellowship; this publication was thus made possible through the support of grant 62192 from the John Templeton Foundation to LSSTC. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of LSSTC or the John Templeton Foundation. R.C. acknowledges support from NASA Swift grant 80NSSC22K0946. The UCSC team is supported in part by NASA grants 80NSSC23K0301 and 80NSSC24K1411; and a fellowship from the David and Lucile Packard Foundation to R.J.F. T.A. acknowledges generous support from the David and Lucile Packard Foundation. C.G. is supported by a VILLUM FONDEN Young Investigator Grant (VIL25501) and Experiment Grant (VIL69896). C.L. acknowledges support under DOE award DE-SC0010008 to Rutgers University. G.N. is funded by NSF CAREER grant AST-2239364, supported in-part by funding from Charles Simonyi. G.N. also gratefully acknowledges NSF support from AST-2206195, OAC-2311355, AST-2432428, as well as AST-2421845 and funding from the Simons Foundation for the NSF-Simons SkAI Institute. G.N. is also supported by the DOE through the Department of Physics at the University of Illinois, Urbana-Champaign (# 13771275), and support from the HST Guest Observer Program through HST-GO-17128 (PI: R. Foley). S.-H.P. was supported by the National Research Foundation of Korea (NRF) NRF-2019R1A2C2010885 and NRF-2022H1D3A2A01096434.
Some/all of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/aj20-5b11 (catalog https://doi.org/10.17909/aj20-5b11). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts.
This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. These observations are associated with programs HST-GO-17497 (PI Valenti) and HST-GO-17772 (PI Bostroem).
This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs JWST-GO-3921 (PI Fox), JWST-DD-4575 (PI Ashall), and JWST-GO-5290 (PI Ashall). The authors acknowledge the teams led by PI Ashall and PI Fox for developing their observing program with a zero-exclusive-access period.
We acknowledge the use of public data from the Swift data archive.
The Young Supernova Experiment (YSE) and its research infrastructure is supported by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (ERC Grant Agreement 101002652, PI K. Mandel), the Heising-Simons Foundation (2018-0913, PI R. Foley; 2018-0911, PI R. Margutti), NASA (NNG17PX03C, PI R. Foley), NSF (AST–1720756, AST–1815935, PI R. Foley; AST–1909796, AST-1944985, PI R. Margutti), the David & Lucille Packard Foundation (PI R. Foley), VILLUM FONDEN (project 16599, PI J. Hjorth), and the Center for AstroPhysical Surveys (CAPS) at the National Center for Supercomputing Applications (NCSA) and the University of Illinois Urbana-Champaign. Pan-STARRS is a project of the Institute for Astronomy of the University of Hawaii, and is supported by the NASA SSO Near Earth Observation Program under grants 80NSSC18K0971, NNX14AM74G, NNX12AR65G, NNX13AQ47G, NNX08AR22G, 80NSSC21K1572, and by the State of Hawaii. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, STScI, NASA under grant NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, NSF grant AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation.
Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the indigenous Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain.
A major upgrade of the Kast spectrograph on the Shane 3 m telescope at Lick Observatory was made possible through generous gifts from the Heising-Simons Foundation as well as William and Marina Kast. Research at Lick Observatory is partially supported by a generous gift from Google.
Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant Nos. AST-1440341, AST-2034437, and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, the University of Washington, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW.
This publication has made use of data collected at Lulin Observatory, partly supported by the TAOvA with the NSTC grant 113-2740-M-008-005.
YSE-PZ was developed by the UC Santa Cruz Transients Team, supported in part by NASA grants NNG17PX03C, 80NSSC19K1386, and 80NSSC20K0953; NSF grants AST–1518052, AST–1815935, and AST–1911206; the Gordon & Betty Moore Foundation; the Heising-Simons Foundation; a fellowship from the David and Lucile Packard Foundation to R. J. Foley; Gordon and Betty Moore Foundation postdoctoral fellowships and a NASA Einstein fellowship, as administered through the NASA Hubble Fellowship program and grant HST-HF2-51462.001, to D. O. Jones; and a National Science Foundation Graduate Research Fellowship, administered through grant No. DGE-1339067, to D. A. Coulter.
Appendix A Additional X-ray Analysis
HEASoft v6.35.1 was used to extract the late-time Swift-XRT spectra. Using xselect v2.5c, we combined observations carried out roughly within a span of two weeks. A 20” source region, and a similar size background region positioned to the northeast of the source, were used to extract source and background spectra using xselect. The small size of the source region was required to exclude any extraneous X-ray sources. Exposure maps were combined using ximage version 4.5.1. The xrtmkarf command was used to create the .arf file, while the swxpc0to12s620210101v016.rmf file was downloaded as the common .rmf calibration file for each observation. The grppha command was then used to prepare the spectra for further analysis. Analysis and fitting of the spectra was done using Sherpa 4.17.0. The minimum column density was set to the Galactic column density obtained from the HI4PI Survey conducted in 2016 ( cm-2). Given the low statistics, we did not group the spectra, but fit both the source and background spectra simultaneously. The Cash statistic was used as the fitting statistic. This is a better statistic for low-counts data where the traditional chi-squared statistic may be unreliable. The tbabs absorption model and apec thermal emission model (xstbabs.abs1*xsvapec.v1) were used to fit the source spectra, while a power law model matched the background spectra reasonably well. In our tbabs model we used the wilm set of abundances (Wilms et al., 2000).
| UT Date | MJD | Phaseaa0.1-30 m | Telescope | Instrument | Wavelength Range | Data Source |
| (days) | (m) | |||||
| 2023-09-06T03:34:00 | 60193.1 | 110.3 | Shane | Kast | 0.33-1.1 | YSE |
| 2023-09-26T02:57:57 | 60213.1 | 130.3 | Shane | Kast | 0.33-1.1 | YSE |
| 2023-11-22T13:31:17 | 60270.6 | 187.7 | Shane | Kast | 0.33-1.1 | YSE |
| 2024-02-14T16:11:48 | 60354.7 | 271.8 | Keck | LRIS | 0.32-1.0 | YSE |
| 2024-03-19T11:38:19 | 60388.5 | 305.7 | Shane | Kast | 0.33-1.1 | YSE |
| 2024-04-12T08:59:13 | 60412.4 | 329.5 | Shane | Kast | 0.33-1.1 | YSE |
| 2024-04-18T08:38:13 | 60418.4 | 335.5 | Shane | Kast | 0.33-1.1 | YSE |
| 2024-04-29T11:10:42 | 60429.5 | 346.63 | Shane | Kast | 0.33-1.1 | YSE |
| 2024-05-17T07:32:49 | 60447.3 | 364.5 | Shane | Kast | 0.33-1.1 | YSE |
| 2024-05-29T07:54:55 | 60459.3 | 376.5 | Shane | Kast | 0.33-1.1 | YSE |
| 2025-03-04T14:20:32 | 60738.6 | 655.8 | Keck | LRIS | 0.32-1.0 | YSE |
| 2023-08-10T05:15:42 | 60166.2 | 83.5 | Shane | Kast | 0.36-1.1 | TReXb |
| 2023-08-10T05:30:49 | 60166.2 | 83.5 | Shane | Kast | 0.58-0.74 | TReXb |
| 2023-08-25T04:14:18 | 60181.2 | 98.4 | Shane | Kast | 0.36-1.1 | TReXb |
| 2023-09-07T03:08:51 | 60194.1 | 111.4 | Shane | Kast | 0.36-1.1 | TReXb |
| 2023-12-05T12:54:31 | 60283.5 | 200.8 | Shane | Kast | 0.36-1.1 | TReXb |
| 2023-12-05T13:12:42 | 60283.6 | 200.9 | Shane | Kast | 0.58-0.74 | TReXb |
| 2024-02-14T10:24:55 | 60354.4 | 271.6 | Shane | Kast | 0.36-1.1 | TReXb |
| 2024-03-18T11:52:24 | 60387.5 | 304.7 | Shane | Kast | 0.36-1.1 | TReXb |
| 2024-03-18T12:40:35 | 60387.5 | 304.7 | Shane | Kast | 0.58-0.74 | TReXb |
| 2024-06-14T06:52:20 | 60475.3 | 392.5 | Shane | Kast | 0.36-1.1 | TReXb |
| 2024-06-14T07:07:56 | 60475.3 | 392.5 | Shane | Kast | 0.58-0.74 | TReXb |
| 2024-08-28T04:06:49 | 60549.2 | 467.4 | Shane | Kast | 0.36-1.1 | TReXb |
| 2024-12-31T15:44:00 | 60675.7 | 592.8 | Keck | KCWI | 0.36-0.89 | Caltech |
| 2025-01-26T15:15:56 | 60701.6 | 618.8 | Keck | LRIS | 0.31-1.0 | Caltech |
| 2025-04-25T12:42:44 | 60790.5 | 707.7 | Keck | LRIS | 0.31-1.0 | Caltech |
| 2024-02-02T15:08:22 | 60342.6 | 259.8 | Keck | NIRES | 0.94-2.5 | KITS (Tinyanont et al., 2024) |
| 2024-02-18T14:56:47 | 60358.6 | 275.8 | Keck | NIRES | 0.94-2.5 | KITS (Tinyanont et al., 2024) |
| 2024-03-01T14:32:38 | 60370.6 | 287.8 | Keck | NIRES | 0.94-2.5 | KITS (Tinyanont et al., 2024) |
| 2024-03-23T14:52:55 | 60392.6 | 309.8 | Keck | NIRES | 0.97-2.5 | KITS (Tinyanont et al., 2024) |
| 2024-03-29T13:43:23 | 60398.6 | 315.7 | Keck | NIRES | 0.97-2.5 | KITS (Tinyanont et al., 2024) |
| 2024-04-26T13:08:20 | 60426.5 | 347.7 | Keck | NIRES | 0.97-2.5 | KITS (Tinyanont et al., 2024) |
| 2024-05-22T11:14:53 | 60452.5 | 369.6 | Keck | NIRES | 1.0-2.5 | KITS (Tinyanont et al., 2024) |
| 2025-03-07T15:14:16 | 60741.6 | 658.8 | Keck | NIRES | 0.94-2.5 | KITS (Tinyanont et al., 2024) |
| 2025-04-12T12:27:10 | 60776.5 | 694.7 | Keck | NIRES | 0.96-2.5 | Caltech |
| 2024-03-24T14:05:30 | 60393.6 | 310.7 | HST | STIS/FUV MAMA | 0.11-0.17 | (Valenti et al., 2024) |
| 2024-03-24T10:55:46 | 60393.5 | 310.6 | HST | STIS/NUV MAMA | 0.16-0.32 | (Valenti et al., 2024) |
| 2025-01-26T17:10:18 | 60393.6 | 619 | HST | STIS/FUV MAMA | 0.11-0.17 | (Bostroem et al., 2024b) |
| 2025-01-26T14:27:10 | 60393.5 | 619 | HST | STIS/NUV MAMA | 0.16-0.32 | (Bostroem et al., 2024b) |
| 2024-01-26T14:34:40 | 60335.6 | 252.8 | JWST | NIRSpec/F170LP | 1.7–3.2 | (Ashall et al., 2023a, b; Medler et al., 2025) |
| 2024-01-26T14:25:02 | 60335.6 | 252.8 | JWST | NIRSpec/F295LP | 2.9–5.3 | (Ashall et al., 2023a, b; Medler et al., 2025) |
| 2024-01-26T15:16:11 | 60335.6 | 252.8 | JWST | MIRI/LRS | 5–14 | (Ashall et al., 2023a, b; Medler et al., 2025) |
| 2024-05-26T13:32:05 | 60456.6 | 373.8 | JWST | NIRSpec/F170LP | 1.7–3.2 | (Ashall et al., 2023a, b; Medler et al., 2025) |
| 2024-05-26T13:17:58 | 60456.6 | 373.8 | JWST | NIRSpec/F295LP | 2.9–5.3 | (Ashall et al., 2023a, b; Medler et al., 2025) |
| 2024-05-26T14:07:53 | 60456.6 | 373.8 | JWST | MIRI/LRS | 5–14 | (Ashall et al., 2023a, b; Medler et al., 2025) |
| 2025-01-08T10:22:19 | 60683.5 | 600.6 | JWST | NIRSpec/F170LP | 1.7–3.2 | (Ashall et al., 2025; Medler et al., 2025) |
| 2025-01-08T10:09:45 | 60683.5 | 600.6 | JWST | NIRSpec/F295LP | 2.9–5.3 | (Ashall et al., 2025; Medler et al., 2025) |
| 2025-01-08T11:02:46 | 60683.5 | 600.6 | JWST | MIRI/LRS | 5–14 | (Ashall et al., 2025; Medler et al., 2025) |
| Phasea | NormFS | NormRS | log(Abs. Flux) | log(Unabs. Flux) | BIC | ||||
| (days) | (keV) | () | ( cm-2) | (keV) | ( cm-2) | (erg s-1 cm-2) | (erg s-1 cm-2) | ||
| 11.1b | – | 0.137b | 5 | 4.7 | |||||
| 11.1b | – | – | – | – | 0.84 | ||||
| 7.0b | – | 0.087b | 2 | 3.6 | |||||
| 7.0b | – | – | – | – | 0.26 | ||||
| 6.5b | – | 0.087b | 1.7 | 4.5 | |||||
| 6.5b | – | – | – | – | 0.70 | ||||